# Quantum geometry and local moment swapover in correlated graphene heterostructures

### v1

@@ Abstract
Graphene-based multilayer systems serve as versatile platforms for exploring the interplay between electron correlation and topology, thanks to distinctive low-energy bands marked by significant quantum metric and Berry curvature from graphene’s Dirac bands. Here, we investigate Mott physics and local spin moments in Dirac bands hybridized with a flat band of localized orbitals in functionalized graphene. Via hybridization control, a topological transition is realized between two symmetry-distinct site-selective Mott states featuring local moments in different Wyckoff positions, with a geometrically enforced metallic state emerging in between. We find that this geometrically controlled local moment “swapover” and associated metal-insulator physics may be realized through proximity coupling of epitaxial graphene on SiC(0001) with group IV intercalants, where the Mott state faces geometrical obstruction in the large-hybridization limit. Our work shows that chemically functionalized graphene provides a correlated electron platform, very similar to the topological heavy fermions in graphene moiré systems but at significantly enhanced characteristic energy scales.
@@
Graphene, initially celebrated as a weakly interacting Dirac material, has developed towards a material basis for correlated electron physics—in multilayer setups with lattice mismatch as well as twisted graphene multilayers, where the moiré potential localizes parts of the electronic states to form flat bands featuring strong electron correlations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. A seemingly complementary platform are moiré-less rhombohedral multilayers like Bernal bilayer, ABC trilayer, and ABCA quadrilayer where correlations emerge in absence of localized electronic orbitals but from itinerant electrons near Van Hove singularities [20, 21, 22, 23, 24, 25, 26, 27]. As a common thread in both platforms, the low-energy states feature sizable Berry curvature and quantum metric, which are inherited from the graphene Dirac cones [28, 29, 30, 31]. These quantum geometric contributions are suggested to play a decisive role in shaping emergent correlated states in these systems [32, 33, 34, 35, 36], including the stabilization of superconducting order [37, 38, 30, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] and various kinds of (pseudo)magnetic states [49, 50, 51, 52, 53, 54, 55].
@@
The emergent ordered states in moiré and rhombohedral multilayers set in at temperature scales of a few Kelvin or below, i.e., $\mathcal{O}$(meV) [20, 2]. What sets these scales is an open matter. Considering the most plausible moiré case from a strong-coupling perspective, the width of the flat bands ($\sim 10$ meV) and the hybridization gap ($\sim 50$ meV) determine order parameter stiffness, therefore also transition temperatures [39, 40, 47, 48]. These energy scales are essentially governed by the graphene interlayer tunellings and moiré potential modulations. It is questionable whether van der Waals engineering can significantly increase them beyond the current stage of graphene multilayer systems.
@@
Figure 1: Flat band hybridized to graphene. (a) Lattice structure of decorated graphene honeycomb lattice with impurity X hybridized to sublattice site A. Only hopping $t$ between sublattices A and B as well as $V$ between X and A exist. (b) Wyckoff positions and their respective local point symmetry groups for the wallpaper group $p3m1$ (No. 156) representing the geometry in panel a. (c) Band structure for different values of $V/t$. The orbital character of the X and graphene atoms (A+B) are colored in red and blue, respectively. (d) Quantum metric $g(\bm{k})=g_{xx}(\bm{k})+g_{yy}(\bm{k})$ of the band crossing the Fermi level at $V/t=0.5$ and $V/t=1.4$. Note the different magnitude of scales. (e) Orbital weight $w_{m}=\sum_{\bm{k}}|w_{\bm{k}m}|^{2}/N_{\bm{k}}$ ($m\in\{\mathrm{A,B,X}\}$) with weight $|w_{\bm{k}m}|^{2}$ of the flat band crossing the Fermi level at $\bm{k}$, and minimal quadratic Wannier function spread $\Omega_{\mathrm{I}}$ as obtained from the quantum metric (c.f. Eq. (2)).
Hybridization control of flat bands: Quantum geometry and localization.  We start with an elemental model, inspired by early works on functionalized graphene [29, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] as well as general considerations for flat band systems like Refs. [69, 70, 71, 72, 73, 74]. We consider a bipartite lattice with sublattices A and B and one atom residing in each sublattice. We will refer to these atoms as A and B atoms, respectively. Sublattice A is furthermore assumed to be decorated with impurity atoms called X atoms. All atom species provide an orbital which transforms trivially under local point group operations. Due to the additional impurity, sublattice sites A and B belong to distinct Wyckoff positions of the crystallographic symmetry group. In Fig. 1a, we illustrate this setup for the example of a decorated two-dimensional honeycomb lattice, i.e., graphene with an impurity X. This geometry belongs to the wallpaper group $p3m1$ (space group 156) with distinct high-symmetry Wyckoff positions 1a, 1b, and 1c on the sublattice sites and hexagon center, respectively, as shown in panel Fig. 1b.
@@
where $\xi(\bm{k})$ is the Fourier transform of the hopping $t$ between A and B atoms and $V>0$ the hopping between the X and the A atoms (c.f. Fig. 1a). For this Hamiltonian at half-filling, a zero-energy eigenstate $|M,\bm{k}\rangle=C_{\bm{k}}(0,V,-\xi(\bm{k}))^{T}$ with $C_{\bm{k}}=1/\sqrt{|\xi(\bm{k})|^{2}+|V|^{2}}$ exists at any $\bm{k}$, independent of additional system specifics. Furthermore, for any finite $V$ there is a finite gap separating the zero-energy band from the two other bands at $\varepsilon_{\pm}(\bm{k})=\pm\sqrt{|\xi(\bm{k})|^{2}+|V|^{2}}$. This is shown for the decorated honeycomb lattice in Fig. 1c where the band structure with orbital/site-weight is shown for different values of $V$.
In the limiting case of dominating hybridization $V\to\infty$, we have $|M,\bm{k}\rangle\to(0,1,0)^{T}$, which means that the flat band’s state is entirely localized at the atoms of the non-decorated sublattice B. This is understandable from a strong bonding-antibonding splitting between the $|X\rangle$ and $|A\rangle$ states that decouple from the system and push all the low-energy spectral weight into sublattice B. Wannieriziation of $|M,\bm{k}\rangle=(0,1,0)^{T}$ leads simply to a Wannier state in real space that is centered and completely peaked in sublattice B, thus the corresponding Wannier spread vanishes.
In the opposite limit of small hybridization $t\gg V\to 0^{+}$, the impurity atoms decouple and constitute the zero energy band $|M,\bm{k}\rangle\to(0,0^{+},-\xi(\bm{k})/|\xi(\bm{k})|)^{T}$ in the entire Brillouin zone except for the nodal point(s) where $\xi(\bm{k})=0$ (c.f. Fig. 1c). The spectral weight of the flat band is correspondingly located at the impurity atoms in this limit. However, since there is no gap closure upon changing the hybridization while keeping it finite ($V>0$), the Wannier center must not have moved away from the B site. Conversely, while the Wannier center remains at B, the maxima of the Wannier centers shift to the three neighboring X sites, spreading over several atoms.
The variation in the Wannier spread is not an artifact but required by the quantum geometry of the zero-energy band: The gauge independent quadratic spread of Wannier functions, $\Omega_{\mathrm{I}}$, is determined by the integral of the quantum metric [75, 76, 77, 37, 78, 32]
@@
which for single-bands measures the spread of maximally localized Wannier functions [77]. The quantum metric $g(\bm{k})$ of the flat band is peaked at the nodes of $\xi(\bm{k})$ for any finite $V>0$. By increasing $V$, it starts to “leak out” by reducing in magnitude and being spread throughout the Brillouin zone, see Fig. 1d. Accordingly, the minimal spread $\Omega_{\mathrm{I}}$ takes a finite value for small $V$ and goes to zero for $V\to\infty$. This is concomitant to the change in X-orbital weight of the flat band as is shown in Fig. 1e. Interestingly, $\Omega_{\mathrm{I}}$ is non-monotonous and displays a maximum for small $V$ despite the quantum metric at the nodes monotonously increasing for $V\to 0^{+}$ ($g(\bm{k}=K)\sim 1/V^{2}$ for the honeycomb lattice in Fig. 1).
Localized orbitals hybridizing with a continuum of Dirac states, where the quantum metric of the resultant flat band is peaked in certain spots of the Brillouin zone, is not only realized in the functionalized graphene cases considered here ($V\to 0^{+}$ limit). It similarly occurs in graphene moiré systems, as highlighted recently by the topological heavy fermion description of magic-angle twisted bilayer graphene [34].
Symmetry-distinct site-selective Mott states.  We now investigate how Coulomb interactions conspire with the quantum geometry for the decorated honeycomb lattice. For simplicity, we assume a local Hubbard interaction $U$ acting on all the atomic species. Our Hamiltonian $\mathcal{H}$ thus has the following new term:
@@
Here, $\uparrow$,$\downarrow$ , ↓ denote the electron spin and $n_{m\uparrow(\downarrow)}$ the number operator. We set the effective local interaction $U=1.6t$ [79, 80], which is well below the critical $U$ for the Mott transition of the undecorated graphene [81]. For vanishing hybridization ($V=0$), graphene is a weakly correlated metal and the flat band of purely X character is a Mott insulator due to the lack of kinetic energy.
The interacting lattice model, $\mathcal{H}=H_{0}+H_{\mathrm{int}}$, is solved within dynamical mean-field theory (DMFT) [82]. The DMFT requires us to solve three independent impurity actions each for A, B, and X sites per self-consistency loop. We solve them using the numerically exact hybridization-expansion continuous-time quantum Monte Carlo method [83, 84]. The self-energy on the Matsubara axis is analytically continued to the real frequency axis using the maximum entropy method [85, 86].
We present the momentum-resolved spectral functions obtained from DMFT in Fig. 2a for three representative regimes of hybridization strength. It exhibits a clear presence of upper and lower Hubbard bands forming Mott gaps ($\sim U$) in both the small and large $V$ regimes [left and right panels] with a flat-band metallic phase intervening for intermediate value of $V$ [middle panel]. Analyzing the underlying self-energy and the spin-spin correlation function reveals that X and B atoms are distinctly responsible for the aforementioned Mottness and concomitant formation of local spin moments in the respective regime. Under particle-hole symmetry, the simplest indicator to this behavior is the imaginary part of the local self-energy $\mathrm{Im}\Sigma(i\omega_{0})$ at the lowest fermionic Matsubara frequency $\omega_{0}=\pi k_{\mathrm{B}}T$ ($T$ is temperature) because $\mathrm{Im}\Sigma$ diverges as $\omega_{n}\to 0$ on the X (B) site for small (large) $V$. Its change with $V$ is shown in Fig. 2b. We note that the emergence of the Dirac cone at the $K$ point for small $V$ in Fig. 2a is attributed to the vanishing weight of the X character on the flat band at the $K$ point ( c.f. Fig. 1b), which is otherwise gapped out everywhere else in the Brillouin zone due to the diverging $\mathrm{Im}\Sigma_{\mathrm{X}}(i\omega_{n}\to 0)$.
To better understand the above behavior, we closely examine the two extreme hybridization limits $V\to 0^{+}$ and $V\to\infty$, where the spectral weight of the non-interacting flat band is predominantly made up of X and B sites, respectively. In these regimes, the local self-energy is well approximated by the Hubbard-I form [87] $\Sigma_{m}(i\omega_{n})=U^{2}/(4i\omega_{n})$ ($m=\mathrm{X}$ at $V\to 0^{+}$ and $m=\mathrm{B}$ at $V\to\infty$). Consequently, $\Sigma_{m}(i\omega_{n})=0$ for the other orbitals $m\in\{\mathrm{A},\mathrm{B}\}$ at $V\to 0^{+}$ and $m\in\{\mathrm{A},\mathrm{X}\}$ at $V\to\infty$. The Hartree term of the self-energy is absorbed into the chemical potential.
@@
Using the analytic form of the Hubbard-I self-energy, the nature of a topological phase transition is revealed by the Luttinger surface—the locus of $\bm{k}$ where the zero-frequency Green’s function $G(\bm{k},\omega=0)$ exhibits zero eigenvalues or, equivalently, where the self-energy diverges at $\omega=0$ [88]. In our case, the Luttinger surface consists of a single band of Green’s function zeros, which is purely of X character at $V\to 0^{+}$ and B character at $V\to\infty$, respectively, because $\mathrm{Im}\Sigma_{m}(i\omega_{n})=-U^{2}/(4\omega_{n})\to-\infty$ for $\omega_{n}\to 0^{+}$ in the respective limit. Since X and B sites reside on distinct Wyckoff positions 1a and 1b (c.f. Fig. 1b), the site-selective Mott states belong to two distinct irreducible representations or two different atomic limits. Hence, the Luttinger surface cannot be deformed continuously by changing $V$ without changing the number of zero bands or breaking symmetries. By preserving all the symmetries as in our DMFT calculations, an intermediate metallic phase as in the middle panel of Fig. 2a has to emerge where the Luttinger surface vanishes. This behavior as a function of $V$ closely resembles the topological phase transition of noninteracting bands in the Su-Schrieffer-Heeger model [89, 90]. In our case, however, it is the topological phase transition of the “Mottness”, i.e., the positions of local spin moments, which is associated with the Green’s function zeros (not poles) forming the Luttinger surface [91, 92].
Figure 3: Hybridization control via proximitity coupling of epitaxial graphene and group-IV triangular adatom lattices. (a, b) Top and side view of the graphene/X/SiC(0001) heterostructure with distance $\Delta z$ between the graphene layer and the X atom buffer layer of group IV elements. The X atoms are in a $\sqrt{3}\times\sqrt{3}$ triangular lattice reconstruction on SiC(0001) and graphene stretched in a $2\times 2$ cell. (c) Energy as a function of layer distance $\Delta z$ for different group IV atoms. The minimum energy $E_{\mathrm{min}}$ is taken as zero and the minimum position is marked with an arrow. In case of X=Ge, two minima close in energy exist. (d) Band structure for the ground state structure for each IV element with the weight of the X-$p_{z}$ and graphene-$p_{z}$ orbitals as fatband plots. For X=Ge, the band structure of the local minimum ($\Delta z=2.7$) is drawn with dashed lines and no orbital weight.
@@
A complimentary route of functionalization is given by proximity coupling of graphene sheets with suitable two-dimensional systems. Here, we focus on epitaxial graphene proximitized to group-IV adatoms at 1/3 monolayer coverage adsorbed on semiconducting surface. Without graphene, these surface lattice systems show various correlated phases [93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. Very recently, it was demonstrated that the addition of graphene to a Sn/SiC(0001) heterostructure host the site-selective Mott states which are strongly governed by hybridization effects [109].
@@
In light of this observation, we investigate the electronic structure of graphene in proximity to different group-IV intercalants (X=C, Si, Ge, Sn, Pb) on a SiC(0001) substrate by using density functional theory (DFT) [110]. The general structure is shown in Figs. 3a and 3b. By tuning the intercalant species X, the distance $\Delta z$ to the graphene layer is strongly changed in the ground state configuration, as the total energy curves in Fig. 3c demonstrate. Here, C intercalants are closest to the graphene layer with the equilibrium distance increasing by over 2 Å going down the periodic table. Notably, the ground state energy for Ge displays two local minima depending on $\Delta z$. The band structure for the lowest-in-energy structure for each X is drawn in Fig. 3d. They all display a relatively flat band pinned to the Fermi level from the intercalant’s $p_{z}$ orbitals hybridized to the Dirac bands of graphene. Due to the variation in equilibrium distances between graphene and the intercalant atoms the hybridization strength is widely tunable.
The band structures obtained from DFT can be rationalized with a generalization of our model to the $2\times 2$ unit cell structure with lower impurity density sketched in Fig. 4a. As in the $1\times 1$ case considered above (c.f. Figs. 1 and 2), we see from Fig. 4b that depending on the hybridization the spectral weight of the flat band moves from the X impurities ($V\to 0^{+}$ limit) to sublattice B ($V\gg t$ limit).
@@
To assess the effect of electron correlations in this generalized model, we perform DMFT calculations using the same parameters as in the $1\times 1$ case. The resultant spectral functions are shown in Fig. 4c. We see that the limit of small $V$ (like $V/t=0.5$) is very similar to the $1\times 1$ case above (c.f. Fig. 2a), as there are well-formed Hubbard bands originating from the X orbitals and a correspondingly large $|\mathrm{Im}\Sigma_{\mathrm{X}}(i\omega_{0})|\gg t$ (c.f. Fig. 4b. Also similar to the $1\times 1$ case, increasing the hybridization quenches the Mott-Hubbard correlations at the X site, as can be seen from the decrease of $|\mathrm{Im}\Sigma_{\mathrm{X}}(i\omega_{0})|$ upon increasing $V$ in Fig. 4b. Likewise, we find a metallic state with a flat band at the Fermi level for intermediate $V/t=1.5$.
@@
The $V\to\infty$ case is interestingly different in the $2\times 2$ case as compared to the $1\times 1$ case. A pseudogap opens with weak but persisting spectral weight of the flat band at the Fermi level even at $V\gg t$ (Fig. 4c, right panel), and we do not find a re-entrant divergent component of the self-energy for $V\gg t$ (Fig. 4d). Instead, the self-energy takes a finite value at several atoms in sublattice B. This behavior can be understood as follows:
Since the density of X atoms is smaller in the $2\times 2$ graphene cell, undecorated A sublattice sites and percolating paths of non-decorated A and B atoms exist, see Fig. 4a. While this does not alter the physics of the small $V$ limit where the spectral weight of the flat band primarily resides on the X atom, it changes the large $V$ limit. Here, the spectral weight becomes distributed over all B sublattice sites and we obtain a robust and finite quantum metric at all $V>0$, see panels b and e of Fig. 4. As a consequence, there is a geometric obstruction preventing the formation of a (site-selective) Mott state in the large $V$ limit. This is captured by the local self-energy in Fig. 4d, which does not diverge on the B sites for $V\to\infty$ in contrast to Fig. 2b. Hence, the momentum-resolved spectral function remains that of a correlated metal with the opening of a pseudogap when the majority of the (non-interacting) flat band’s spectral weight lies on the B site, see Fig 4c. We added in Fig. 4d approximate hybridization values obtained from fitting tight-binding models to the DFT band structures from Fig. 3d, demonstrating that both hybridization limits are captured throughout all group IV atoms.
The correlated metallic state emerging here in the large $V$ regime is stabilized by the quantum geometry of the flat band. The $V\to\infty$ limit for this band is very similar to the vacancies and the vacancy induced flat band of Ref. [31].

### v3

# Real-Space Switching of Local Moments Driven by Quantum Geometry in Correlated Graphene Heterostructures

@@ Abstract
Graphene-based multilayer systems serve as versatile platforms for exploring the interplay between electron correlation and topology, thanks to distinctive low-energy bands marked by significant quantum metric and Berry curvature from graphene’s Dirac bands. Here, we investigate Mott physics and local spin moments in Dirac bands hybridized with a flat band of localized orbitals in functionalized graphene. Via hybridization control, a topological transition is realized between two symmetry-distinct site-selective Mott states featuring local moments in different Wyckoff positions, with a geometrically enforced metallic state emerging in between. We find that this geometrically controlled real-space switching of local moments and associated metal-insulator physics may be realized through proximity coupling of epitaxial graphene on SiC(0001) with group IV intercalants, where the Mott state faces geometrical obstruction in the large-hybridization limit. Our work shows that chemically functionalized graphene provides a correlated electron platform, very similar to the topological heavy fermions in graphene moiré systems but at significantly enhanced characteristic energy scales.
@@
Graphene, initially celebrated as a weakly interacting Dirac material, has developed towards a material basis for correlated electron physics—particularly in twisted graphene multilayers, where the moiré potential localizes parts of the electronic states to form flat bands featuring strong electron correlations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. A seemingly complementary platform are moiré-less rhombohedral multilayers like Bernal bilayer, ABC trilayer, and ABCA tetralayer where correlations emerge in absence of localized electronic orbitals but from itinerant electrons near Van Hove singularities [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. As a common thread in both platforms, the low-energy states feature sizable Berry curvature and quantum metric, which are inherited from the graphene Dirac cones [32, 33, 34, 35]. These quantum geometric contributions are suggested to play a decisive role in shaping emergent correlated states in these systems [36, 37, 38, 39, 40], including the stabilization of superconducting order [41, 42, 34, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] and various kinds of (pseudo)magnetic states [53, 54, 55, 56, 57, 58, 59].
@@
The emergent ordered states in moiré and rhombohedral multilayers set in at temperature scales of a few Kelvin or below, i.e., $\mathcal{O}$(meV) [20, 2]. What sets these scales is an open matter. Considering the most plausible scenario for the moiré case from a strong-coupling perspective, the width of the flat bands ($\sim 10$ meV) and the hybridization gap ($\sim 50$ meV) determine order parameter stiffness, therefore also transition temperatures [43, 44, 51, 52]. These energy scales are essentially governed by the graphene interlayer tunellings and moiré potential modulations. It is questionable whether van der Waals engineering can significantly increase them.
@@
Figure 1: Flat band hybridized to graphene. (a) Lattice structure of decorated graphene honeycomb lattice with impurity X hybridized to sublattice site A created with VESTA [60]. Only hopping $t$ between sublattices A and B as well as $V$ between X and A exist. (b) Wyckoff positions and their respective local point symmetry groups for the wallpaper group $p3m1$ (No. 156) representing the geometry in panel a. (c) Orbital weight $w_{m}=\sum_{\bm{k}}|w_{\bm{k}m}|^{2}/N_{\bm{k}}$ ($m\in\{\mathrm{A,B,X}\}$) with weight $|w_{\bm{k}m}|^{2}$ of the flat band crossing the Fermi level at $\bm{k}$, and minimal quadratic Wannier function spread $\Omega_{\mathrm{I}}$ as obtained from the quantum metric (c.f. Eq. (2) and panel e). (d) Band structure for different values of $V/t$. The orbital character of the X and graphene atoms (A+B) are colored in red and blue, respectively. (e) Quantum metric $\operatorname{\mathrm{Tr}}g(\bm{k})=g_{xx}(\bm{k})+g_{yy}(\bm{k})$ of the band crossing the Fermi level for $V/t=0.5$ (left) and $V/t=3.0$ (right). Note the different magnitude of scales. The middle panels show corresponding Wannier functions which change the maxima from the X sites ($V/t=0.5$) to the B sites ($V/t=3.0$) by increasing $V$.
Hybridization control of flat bands: Quantum geometry and localization.  We start with an elemental model, inspired by early works on functionalized graphene [33, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] as well as general considerations for flat band systems like Refs. [74, 75, 76, 77, 78, 79]. We consider a bipartite lattice with sublattices A and B and one atom residing in each sublattice. Furthermore, sublattice A is assumed to be decorated with impurity atoms called X atoms. All atom species provide an orbital which transforms trivially under local point group operations. Due to the additional impurity, sublattice sites A and B belong to distinct Wyckoff positions of the crystallographic symmetry group. In Fig. 1a, we illustrate this setup for the example of a decorated two-dimensional honeycomb lattice, i.e., graphene with an impurity X. This geometry belongs to the wallpaper group $p3m1$ (space group 156) with distinct high-symmetry Wyckoff positions 1a, 1b, and 1c on the sublattice sites and hexagon center, respectively, as shown in panel Fig. 1b.
@@
where $\xi(\bm{k})$ is the Fourier transform of the hopping $t$ between A and B atoms and $V>0$ the hopping between the X and the A atoms (c.f. Fig. 1a). For this Hamiltonian at half-filling, a zero-energy eigenstate $|M,\bm{k}\rangle=C_{\bm{k}}(0,V,-\xi(\bm{k}))^{T}$ with $C_{\bm{k}}=1/\sqrt{|\xi(\bm{k})|^{2}+|V|^{2}}$ exists at any $\bm{k}$, independent of additional system specifics. Only X and B sites contribute to this flat band with the orbital weight being tuned by the hybridization. Furthermore, for any finite $V$ there is a finite gap separating the zero-energy band from the two other bands at $\varepsilon_{\pm}(\bm{k})=\pm\sqrt{|\xi(\bm{k})|^{2}+|V|^{2}}$. This is shown for the decorated honeycomb lattice in Figs. 1c and 1d where the band structure with orbital/site-weight is shown for different values of $V$.
In the limit of dominating hybridization $V\to\infty$, the flat band’s eigenstate becomes $|M,\bm{k}\rangle\to(0,1,0)^{T}$, i.e., it is entirely localized on the non-decorated B sublattice. This is understandable from a strong bonding-antibonding splitting between $|\mathrm{X}\rangle$ and $|\mathrm{A}\rangle$ states that decouple from the system and push all the low-energy spectral weight into sublattice B. The corresponding Wannier function of $|M,\bm{k}\rangle=(0,1,0)^{T}$ is centered and sharply peaked on the B site, thus yielding a vanishing Wannier spread.
In the opposite limit $t\gg V\to 0^{+}$, the impurity atoms decouple and constitute the zero-energy band $|M,\bm{k}\rangle\to(0,0^{+},-\xi(\bm{k})/|\xi(\bm{k})|)^{T}$ in the entire Brillouin zone except for the nodal point(s) where $\xi(\bm{k})=0$ (c.f. Fig. 1d). The flat-band weight is correspondingly located at the impurity atoms. However, since there is no gap closure upon changing the hybridization while keeping it finite ($V>0$), the Wannier center must not have moved away from the B site. Conversely, while the Wannier center remains at B, the maxima shift to the three neighboring X sites, spreading over several atoms (c.f. insets in Fig. 1e).
The variation in the Wannier spread is not an artifact but required by the quantum geometry of the zero-energy band: The gauge independent quadratic spread of Wannier functions, $\Omega_{\mathrm{I}}$, is determined by the integral of the quantum metric [80, 81, 82, 41, 83, 36, 84]
@@
which for single-bands measures the spread of maximally localized Wannier functions [82]. Here, the quantum metric $\operatorname{\mathrm{Tr}}g(\bm{k})=g_{xx}(\bm{k})+g_{yy}(\bm{k})$ [84] of the flat band is peaked at the nodes of $\xi(\bm{k})$ for any finite $V>0$. By increasing $V$, it “leaks out” by reducing in magnitude and being spread throughout the Brillouin zone, see Fig. 1e. Accordingly, the minimal spread $\Omega_{\mathrm{I}}$ decreases with increasing $V$ and vanishes for $V\to\infty$, concomitant to the reduction in X-orbital weight in the flat band as shown in Fig. 1c. Interestingly, $\Omega_{\mathrm{I}}$ is non-monotonic and displays a maximum for small $V$, despite the quantum metric at the nodes increasing monotonically as $V\to 0^{+}$ ($g(\bm{k}=\bm{K})\sim 1/V^{2}$ for the honeycomb lattice in Fig. 1 [84]).
This structure of localized orbitals hybridizing with a continuum of Dirac states and producing flat bands with a quantum metric peaked at specific spots in the Brillouin zone is not unique to functionalized graphene. It similarly occurs in graphene moiré systems, as highlighted by the topological heavy fermion description of magic-angle twisted bilayer graphene [38, 111].
Symmetry-distinct site-selective Mott states.  We now investigate how Coulomb interactions conspire with the quantum geometry for the decorated honeycomb lattice. We assume a local Hubbard interaction $U$ acting on all the atomic species:
@@
Here, $\uparrow,\downarrow$ denote the electron spin and $n_{m\uparrow(\downarrow)}$ the number operator. We set the effective local interaction $U=1.6t$ [106, 107], which is well below the critical $U$ for the Mott transition of the undecorated graphene [112]. For vanishing hybridization ($V=0$), graphene is a weakly correlated metal and the flat band of purely X character is a Mott insulator due to the lack of kinetic energy.
The interacting lattice model, $\mathcal{H}=H_{0}+H_{\mathrm{int}}$, is solved within dynamical mean-field theory (DMFT) [113]. The DMFT requires us to solve three independent impurity actions each for A, B, and X sites per self-consistency loop. We solve them using the numerically exact hybridization-expansion continuous-time quantum Monte Carlo method [114, 115]. The self-energy on the Matsubara axis is analytically continued to real frequencies using the maximum entropy method [116, 117].
We present the momentum-resolved spectral functions obtained from DMFT in Fig. 2a for three representative regimes of hybridization strength. In both the small and large $V$ regime [left and right panels], the spectra exhibit clear upper and lower Hubbard bands forming Mott gaps ($\sim U$) with a flat-band metallic phase intervening at intermediate $V$ [middle panel]. Analyzing the underlying self-energy and spin-spin correlation function [84] reveals that X and B atoms distinctly drive the aforementioned Mottness and concomitant formation of local spin moments in the small and large $V$ regimes, respectively. Under particle-hole symmetry, this is captured by the imaginary part of the local self-energy $\mathrm{Im}\,\Sigma(i\omega_{0})$ at the lowest fermionic Matsubara frequency $\omega_{0}=\pi k_{\mathrm{B}}T$ ($T$ is temperature), because $\mathrm{Im}\Sigma$ diverges as $\omega_{n}\to 0$ on the X (B) site for small (large) $V$ as shown in Fig. 2b. We note that the Dirac cone at $K$ for small $V$ arises from the vanishing X character of the flat band at this point (c.f. Fig. 1b), which is otherwise gapped out everywhere else in the Brillouin zone due to the diverging $\mathrm{Im}\Sigma_{\mathrm{X}}(i\omega_{n}\to 0)$
To understand this behavior, we focus on the extreme limits $V\to 0^{+}$ and $V\to\infty$, where the spectral weight of the non-interacting flat band predominantly resides on X and B sites, respectively. In these limits, the local self-energy is well approximated by the Hubbard-I form [118] $\Sigma_{m}(i\omega_{n})=U^{2}/(4i\omega_{n})$ for $m=\mathrm{X}$ at $V\to 0^{+}$ and $m=\mathrm{B}$ at $V\to\infty$, while $\Sigma$ vanishes on the other orbitals. The Hartree term is absorbed into the chemical potential.
@@
Using the analytic form of the Hubbard-I self-energy, the nature of a topological phase transition is revealed by the Luttinger surface—the locus of $\bm{k}$ where the zero-frequency Green’s function $G(\bm{k},\omega=0)$ exhibits zero eigenvalues or, equivalently, where the self-energy diverges at $\omega=0$ [91]. Here, the Luttinger surface consists of a single band of Green’s function zeros, which is purely of X character at $V\to 0^{+}$ and of B character at $V\to\infty$ [84]. Since X and B occupy distinct Wyckoff positions (c.f. Fig. 1b), the site-selective Mott states belong to distinct irreducible representations, i.e., different atomic limits. Hence, the Luttinger surface cannot be deformed continuously by changing $V$ without changing the number of zero bands or breaking symmetries. Consequently, an intermediate metallic phase as in the middle panel of Fig. 2a has to emerge where the Luttinger surface vanishes. This behavior closely resembles the topological phase transition of noninteracting bands in the Su-Schrieffer-Heeger model [119, 120]. In our case, however, it is the topological phase transition of the “Mottness”, i.e., the positions of local spin moments, which is associated with the Green’s function zeros (not poles) forming the Luttinger surface [121, 122, 123, 124, 125, 92, 126, 127, 128, 129, 130, 131]. We note that these observations are robust against longer-range hoppings or on-site terms [84].
@@
A complementary route of functionalization is given by proximity coupling of graphene sheets with suitable two-dimensional systems. Here, we focus on epitaxial graphene proximitized to group-IV adatoms at 1/3 monolayer coverage adsorbed on semiconducting surface. Without graphene, these surface lattice systems show various correlated phases [132, 133, 134, 135, 136, 137, 138, 139, 104, 105, 140, 141, 142, 143, 144, 145, 100].
@@
Motivated by this observation, we investigate the electronic structure of graphene in proximity to different group-IV intercalants (X=\,=\,C, Si, Ge, Sn, Pb) on a SiC(0001) substrate by using density functional theory (DFT) [84]. The structural details are discussed in the End Matter. The band structure for for each IV element X is drawn in Fig. 3. They all display a relatively flat band pinned to the Fermi level from the intercalant’s $p_{z}$ orbitals hybridized to the Dirac bands of graphene. Due to the variation in equilibrium distances between graphene and the intercalant atoms, the hybridization strength and orbital character is widely tunable. The site- and orbital-resolved density of states reveals that dominantly X-$p_{z}$ and B-site graphene-$p_{z}$ contribute to the formation of the flat band [84].
[…]
@@
The band structures obtained from DFT can be rationalized by extending model (1) to a $2\times 2$ unit cell structure with reduced impurity density sketched in Fig. 4a; model details are given in the End Matter. The resulting spectral functions are shown in Fig. 4c. The limit of small $V$ ($V/t=0.5$) is very similar to the $1\times 1$ case (c.f. Fig. 2a), showing well-formed Hubbard bands originating from the X orbital and a correspondingly large $|\mathrm{Im}\,\Sigma_{\mathrm{X}}(i\omega_{0})|\gg t$ (c.f. Fig. 4b). Increasing the hybridization quenches Mott-Hubbard correlations at the X site, as indicated by the reduction of $|\mathrm{Im}\,\Sigma_{\mathrm{X}}(i\omega_{0})|$ upon increasing $V$ in Fig. 4b. Likewise, we find a metallic state with a flat band at the Fermi level for intermediate $V/t=1.5$.
@@
The $V\to\infty$ case is interestingly different in the $2\times 2$ case as compared to the $1\times 1$ case. A pseudogap opens with weak but persisting spectral weight of the flat band at the Fermi level even at $V\gg t$ (Fig. 4c, right panel), and we do not find a re-entrant divergent component of the self-energy for $V\gg t$ (Fig. 4b). Instead, the self-energy takes a finite value on several atoms in sublattice B [84].
This qualitative difference originates from the lower density of X sites in the $2\times 2$ lattice, as it allows for undecorated A sites and percolating A–B paths (c.f. Fig. 4a). While the small-$V$ regime is still dominated by X-site physics, at large $V$ the flat-band spectral weight redistributes over the B sublattice, resulting in a finite quantum metric for all $V>0$ (see End Matter). This geometric obstruction prevents the formation of a (site-selective) Mott state in the strong-hybridization limit. Accordingly, the local self-energy remains finite on the B sites and the system stays a correlated metal with a pseudogap when the flat-band weight predominantly resides on those sites (Fig. 4c). The hybridization values relevant for the DFT band structures of Fig. 3, extracted from tight-binding fits [84], span all regimes and are indicated in Fig. 4b.